OpenStudy (jonspiro):

A function has f(x) has a limit L at x = a if and only if it has both left and right limits there and these one­-sided limits are both equal to L. That is; limx→a f(x) = L ⇔ limx→a f(x) = limx→a+ f(x) = L . Use this fact about the relationship between one­sided limits and two­sided limits to answer the following questions a) If f(x)=|x2| / (x^2+x-6) ,find i. limx→2+ f(x), ii. limx→2- f (x), iii. limx→2 f (x), b) What can you say about the relationship between the solved functions in (a)?

OpenStudy (irishboy123):

\(\lim\limits_{x \to 2} \dfrac{|x|}{x^2 + x - 6}\) \(= 2 \lim\limits_{x \to 2} \dfrac{1}{x^2 + x - 6}\) problem is that the denom goes to zero we can investigate the left and right-sidedness using a calculator ... or by saying that we let \(x = 2 + \delta\), so \( x- 2 = \delta \), where \( \delta <<1\) and the limit is now \(\delta \to 0\) \(= 2 \lim\limits_{\delta \to 0} \dfrac{1}{(2+\delta)^2 + 2+ \delta - 6}\) \(= 2 \lim\limits_{\delta \to 0} \dfrac{1}{5 \delta }\)....... ignoring \(\delta^2\) terms in the denom You can see that we still get the singularity but that it is \(- \infty \) where \(\delta < 0\), ie the left-sided limit; and \(+ \infty \) where \(\delta > 0\), ie the right-sided limit IOW \(\lim\limits_{x \to 2^-} \dfrac{|x|}{x^2 + x - 6} = - \infty\) \(\lim\limits_{x \to 2^+} \dfrac{|x|}{x^2 + x - 6} = \infty\) So there is no limit or derivative at \(x = 2\)

satellite73 (satellite73):

it is \(|x^2|\) up top, right?

OpenStudy (irishboy123):

does that matter ?

OpenStudy (jonspiro):

oops my bad it is |x-2| on top

OpenStudy (jonspiro):

If f(x)=|x-2| / (x^2+x-6) ,find i. limx→2+ f(x), ii. limx→2- f (x), iii. limx→2 f (x),

satellite73 (satellite73):

oooh \(|x-2|\) !!

OpenStudy (jonspiro):

yes

satellite73 (satellite73):

if \(x>2\) then \(|x-2|=x-2\) if \(x<2\) then \(|x-2|=2-x\)

OpenStudy (irishboy123):

so it's indeterminate good spot satellite you can factor the bottom: \[\lim\limits_{x \to 2} \dfrac{|x-2|}{(x-2)(x+3)}\]

satellite73 (satellite73):

so the right handed limit will be for \[\frac{x-2}{x^2+x-6}\] and the left handed lmit will be for \[\frac{2-x}{x^2+x-6}\]

satellite73 (satellite73):

factor and cancel, see when you get for the limit should be pretty clear that one will be the negative of the other

OpenStudy (irishboy123):

if \(x<2\) you have \[\lim\limits_{x \to 2^-} \dfrac{-(x-2)}{(x-2)(x+3)}\] if \(x>2\) you have \[\lim\limits_{x \to 2^+} \dfrac{x-2}{(x-2)(x+3)}\] and cos we are playing with limits and never actually getting to 2, you can do the algebra \[\lim\limits_{x \to 2^-} \dfrac{-(x-2)}{(x-2)(x+3)} \to -\dfrac{1}{2+3}\] \[\lim\limits_{x \to 2^+} \dfrac{x-2}{(x-2)(x+3)} \to \dfrac{1}{2+3}\]

OpenStudy (jonspiro):

can i post another question on here?

OpenStudy (irishboy123):

opening a new thread, and closing this one, is usually best advice :)

OpenStudy (jonspiro):

ok

OpenStudy (jonspiro):

relationship between the solved functions? any comments on that?

OpenStudy (irishboy123):

don't actually understand the question. seems to me there is only 1 function :(